# Probing a light long-lived pseudo-scalar from Higgs decay via displaced taus at the LHC

### v1

@@ 1 Introduction
Theoretically, a light long-lived pseudo-scalar can be readily predicted in some specific 2HDMs or in some non-minimal supersymmetric models like the next-to-minimal supersymmetric model  Ellwanger:2009dp ; Cao:2012fz ; Cao:2013gba ; Cao:2010na ; Adhikary:2024dix . For example, models like 2HDM+$a$ can provide a very light pseudo scalar $a$ as LLP and has been widely studied Haisch:2023rqs . And in the type-I 2HDM, where the couplings of the pseudo-scalar $A$ with fermions are proportional to $1/\tan\beta$, a very light $A$ can naturally act as a LLP, as studied for the experiments like FASER and MATHUSLA  Kling:2022uzy ; Liu:2024azc . In this work, we take the aligned 2HDM as an example and focus on a long-lived pseudo-scalar $A$ produced from the Higgs decay $h\to AZ$ at the LHC. Different from the type-I 2HDM, the Yukawa couplings of $A$ in the aligned 2HDM are proportional to three different factors: $\kappa_{\ell}$ for leptons, $\kappa_{d}$ for down-type quarks, and $\kappa_{u}$ for up-type quarks. Assuming $\kappa_{d}=\kappa_{u}=0$, a light $A$ decays dominantly to $\tau^{+}\tau^{-}$, and it can become a LLP for a sufficient small $\kappa_{\ell}$. This light long-lived $A$ can be produced from the Higgs decay $h\to AZ$ at the LHC.
This long-lived pseudo-scalar can fly over dozens to hundreds of millimeters once produced in the Higgs decay $h\to AZ$ at the LHC. The tau leptons from its decay ($A\to\tau^{+}\tau^{-}$) will further decay into tracks which can be reconstructed only if they traverse sufficient layers in an Inner Detector (Tracker). This has been implemented in the ATLAS and CMS experiments as a special technique of tracking ATLAS:2023nze ; CMS:2014pgm . Upon these Large Radius Tracks (LRT), it becomes even possible to reconstruct the decaying vertex of the LLP ATLAS:2019wqx ; CMS:2021tkn . In this work we will use these new techniques as performed in  ATLAS:2020xyo . On the other hand, the tau jet with a displaced track provides an outstanding signature to our signal. Since the light pseudo-scalar is produced from the decay of a not-so-heavy Higgs boson ($\sim 125$ GeV ) in association with a rather heavy $Z$-boson ($\sim 91$ GeV), its momentum is usually not large, such that its decay products are usually soft. Considering the softness of the taus from the pseudo-scalar decay, in this work we propose to reconstruct a merged (joint) vertex from the decays of the two taus in order to suppress backgrounds. Although this reconstruction has larger uncertainties than the conventional (standard) vertexing, as will be demonstrated in our study, it can help to discriminate our signal from backgrounds.
@@ 2.2 Current theoretical and experimental constraints
In our scan $m_{12}^{2}$ is also a free parameter whose value is tuned to satisfy the theoretical constraints from the vacuum stability and perturbativity as well as unitarity. For the Yukawa sector, we set $\kappa_{u}=\kappa_{d}$ to zero and scan $\kappa_{\ell}$ within a range of small values
@@
    The constraints by the signal strength of the 125 GeV Higgs. The non-vanishing $\cos(\beta-\alpha)$ in the couplings of $h$ to the SM particles, is constrained by HiggsSignals  Bechtle:2013xfa package to satisfy all the observations (within $2\sigma$). The decay $h\to AA$ is bounded by the invisible Higgs decay of $BR(h\to{\rm invisible})<0.24$  CMS:2016dhk ; ATLAS:2015gvj ; ATLAS:2017nyv .
@@
In Fig.1 we present the scatter plots of the samples that satisfy the aforementioned constraints. The following characteristics can be observed from this figure:
@@
    The top-left panel shows that $\cos(\beta-\alpha)$ is constrained to be below 0.07 for $m_{A}=5{\rm~{}GeV}$ and below 0.18 for $m_{A}=30{\rm~{}GeV}$. The most stringent constraints on $\cos(\beta-\alpha)$ arise from the measurements of the Higgs boson discovered at 125 GeV.
@@
    The top-right panel shows that $\tan\beta$ is favored in the range 0.6 and 1.6, as explained in Appendix C.
@@
    The bottom-left panel indicates that the invariant lifetime $c\tau$ of $A$ is primarily determined by $\kappa_{\ell}$ as expected and is inversely proportional to $\kappa_{\ell}^{2}$.
  • (4)
    The bottom-right panel shows that the branching ratio of $h$ decay to $A$ and $Z$, i.e $BR(h\to ZA)$, is proportional to $\cos^{2}(\beta-\alpha)$, with a maximum value of approximately 0.068. Given that the cross-section for Higgs production via gluon fusion equals $48.58$ pb LHCHiggsCrossSectionWorkingGroup:2016ypw , the maximum cross section for $gg\to h\to ZA$ is approximately 3.3 pb.
@@ 3 Observability at the LHC
To save computing time in numerical computation, only a few representative benchmark points are chosen from the allowed parameter space for an illustration, as listed in Table 1. We choose low mass at 5 GeV and high mass at 30 GeV, and vary the invariant lifetime from 3 mm to 300 mm. We use FeynRules Alloul:2013bka to generate the UFO model file, MadGraph5_aMC@NLO Alwall:2014hca ; Frederix:2018nkq to generate parton-level events, Pythia8 Bierlich:2022pfr ; Sjostrand:2014zea for parton showering and hadronization, and then prepare signal events in HepMC3 format. The gluon-gluon fusion Higgs production cross section is scaled to the N3LO value LHCHiggsCrossSectionWorkingGroup:2016ypw . For the $t\bar{t}$ background, we require the top quark to decay into a $W$-boson and a $b$ quark, with the $W$-boson subsequently decaying leptonically, while the $t\bar{t}$ cross section is scaled to the NNLO+NNLL value Czakon:2011xx . Among the $Z\tau\tau$ backgrounds, the $Z\gamma^{\ast}$ component is estimated to be about $38\%$ by MadGraph5_aMC@NLO. As a crosscheck, we show the distributions of the radius of particle production position in Fig.2, which demonstrates that these samples are generated according to physics expect.
Table 1: Physics parameters for benchmark points in the numerical study. The $tan\beta$ is taken at $0.8$ ($1.3$) for low (high) masses, $m^{2}_{12}$ is taken at $3900$ ($4200$) GeV2 for low (high) masses. The value of $\cos(\beta-\alpha)$ is 0.03 for all points.
@@ 3.1 The fast simulation and selection procedure
The framework Delphes deFavereau:2013fsa is adopted in this work. The setup of detector responses mainly follows that of ATLAS with pileup therein. New modules are developed as described below to accommodate a long-lived particle, in consistency with the methodology described in ATLAS:2023nze ; CMS:2022prd ; atlas:2019pn ; ATLAS:2019wqx ; CMS:2021tkn  [^3] These algorithms and the response parameters therein are followed just for a quick estimation of physics potential.:
@@
    Displaced tracks. In the scenario of LLP with a lifetime around dozens of millimeters, tracks from the LLP decay will be produced within the Tracker but somewhat far away from the collision point. Typically this can be reflected by the transverse impact parameter $d_{0}$ as shown in Fig.3. In our tracking module of Delphes, $d_{0}\geq 10~{}mm$ is adopted to define the displaced tracks, while other setup as in ATLAS:2023nze . There is also an observation that the mass difference around 35 GeV $(\simeq m_{h}-m_{Z})$ leaves only less energy to tracks, so a low threshold $p_{T}\geq 0.5$ GeV is used for the track transverse momentum. But a track produced too far ($R_{v}\geq 300~{}mm$) will leave too few effective hits within the Inner Detector and thus will be rejected. From Figs.2 and 3 we see that with a fixed $c\tau$ value a lighter pseudo-scalar is more likely to travel a longer distance because it tends to have a higher velocity.
@@
    Displaced $\tau$ jet ($\tau$ jet with displaced tracks). A tau jet basing on the performance described in atlas:2019pn is expected to be associated with one or three extra displaced tracks within a cone $\sqrt{{(\Delta\phi)}^{2}+{(\Delta\eta)}^{2}}\leq 0.2$, as aforementioned. To keep as many tau signal events as possible, the threshold of tau $p_{T}$ is reduced to 10 GeV from the conventional cut like 20 GeV. For the taus with $p_{T}$ below 20 GeV, an efficiency degradation by $60\%$ is supposed as a compensation. Meanwhile, the possibility of a $b$-jet mis-identified as a tau jet is adjusted from $5\%$ to $10\%$ for $p_{T}$ within 10-20 GeV. For the events with only one hadronic tau tagged, a displaced muon is required as a substitution.
@@
    Reconstruction of $Z$-boson mass and visible mass of Higgs boson. As usual, the invariant mass for a pair of same-flavor and opposite-sign leptons is required to fall in a window of $81{\rm~{}GeV}<m_{\ell\ell}<101{\rm~{}GeV}$, with isolated leptons in the phase space of $p_{T}^{\ell}>15{\rm~{}GeV}$ and $|\eta|<2.5$. No doubt this $Z$-peak (in fact the leptons’ isolation) can suppress the QCD backgrounds down to a sufficiently low level. Also, by adding up the four-momenta of a pair of charged leptons (from $Z$) and a pair of tau candidates, a visible mass is reconstructed to suppress events other than the Higgs decay. As shown in Fig.5, it helps to reject all backgrounds.
@@
Table 2: The separately estimated efficiencies for the backgrounds and the signals at each benchmark points. The cross-section of each background prior to cuts is shown in square brackets in unit of pb.
$m_{Z}$-window displaced $\tau$-jet displaced vertex $m^{vis}_{h}\leq 125$ GeV final efficiency
$Z\tau\tau~{}[0.027]$ 0.4 $1.4\times{10}^{-3}$ $7.9\times{10}^{-2}$ $7.5\times{10}^{-3}$ $3.3\times{10}^{-7}$
$Zbb~{}[68.2]$ 0.4 $9.8\times{10}^{-5}$ $6.9\times{10}^{-2}$ $9.9\times{10}^{-2}$ $4.1\times{10}^{-7}$
$t{\bar{t}}~{}[92.4]$ $2.2\times{10}^{-2}$ $2.5\times{10}^{-4}$ $9.1\times{10}^{-2}$ $2.5\times{10}^{-2}$ $1.7\times{10}^{-8}$
BP1 0.44 $3.2\times{10}^{-3}$ $8.2\times{10}^{-2}$ 0.95 $0.051\%$
@@
The performance of these simulation procedures is summarized in Table 2. For illustration, the efficiency is shown independently at each step. In fact, there are correlations among these cuts, especially between displaced tau-jet and displaced vertex, so the final efficiency differs from the simple product of these steps. As shown, the signal events can be selected at a few thousandths, while the backgrounds are about 4 or 5 orders lower. For the low (high) mass of 5 (30) GeV, the efficiency reaches the maximal value when $c\tau$ is dozens (hundreds) of millimeters, which indicates a lighter pseudo-scalar travels faster. Generally, the high mass of 30 GeV gives a higher efficiency than the low mass of 5 GeV because a light pseudo-scalar gives soft decay products. Obviously, from the two groups of BP1-BP4 and BP5-BP8 we see a very short (long) lifetime gives a low efficiency because the pseudo-scalar cannot reach (go beyond) the sensitive zones of the detector. During these selection procedures, we find that the pileup entered by contaminating displaced tracks, playing a dominant role in backgrounds.
@@ 3.2 Results and discussions
Figure 6: The red and orange denote signal event yields with $BR(h\to ZA)\simeq{10}^{-3}$, with the red for $(m_{A}=30{\rm~{}GeV},~{}c\tau=30{\rm~{}mm})$ and orange for $(m_{A}=5{\rm~{}GeV},~{}c\tau=30{\rm~{}mm})$. Other colors denote the backgrounds.
Table 3: The first row is the expected limit on $BR(h\to ZA)$ estimated with pure backgrounds, while the second row is the expected $BR(h\to ZA)$ to reach the expected $3\sigma$ significance through the selection procedures, all numbers are in percents.
BP1 BP2 BP3 BP4 BP5 BP6 BP7 BP8
limit 2.57 0.87 1.20 3.74 0.41 0.29 0.25 0.35
at $3\sigma$ 12 3.9 5.4 17 2.3 1.6 1.3 2.0

### v2

@@ 1 Introduction
Theoretically, a light long-lived pseudo-scalar can be readily predicted in some specific 2HDMs or in some non-minimal supersymmetric models like the next-to-minimal supersymmetric model  Ellwanger:2009dp ; Cao:2012fz ; Cao:2013gba ; Cao:2010na ; Adhikary:2024dix . For example, models like 2HDM+$a$ can provide a very light pseudo scalar $a$ as LLP and has been widely studied Haisch:2023rqs . Alternatively in a type-I 2HDM, where the couplings of the pseudo-scalar $A$ with fermions are proportional to $1/\tan\beta$, a very light $A$ can naturally act as a LLP, as studied for the experiments like FASER and MATHUSLA  Kling:2022uzy ; Liu:2024azc . In this work, we take however an aligned 2HDM as an example to study a long-lived pseudo-scalar $A$ produced from the Higgs decay $h\to AZ$ at the LHC. In this aligned 2HDM the Yukawa couplings of $A$, different from the type-I 2HDM, are proportional to three different factors: $\kappa_{\ell}$ for leptons, $\kappa_{d}$ for down-type quarks, and $\kappa_{u}$ for up-type quarks, so $A$ can become a LLP when all Yukawa couplings are sufficiently small. In this work we investigate the decay of $A$ to $\tau^{+}\tau^{-}$ for its larger yields (relative to other leptons) and better rejection against backgrounds in experiment (compared with decay to quarks).
This long-lived pseudo-scalar can fly over dozens to hundreds of millimeters once produced in the Higgs decay $h\to AZ$ at the LHC. The tau leptons from its decay ($A\to\tau^{+}\tau^{-}$) will further decay into tracks which can be reconstructed only if they traverse sufficient layers in an Inner Detector (Tracker). This has been implemented in the ATLAS and CMS experiments as a special technique of tracking ATLAS:2023nze ; CMS:2014pgm . Upon these Large Radius Tracks (LRT), it becomes even possible to reconstruct the decaying vertex of the LLP ATLAS:2019wqx ; CMS:2021tkn . In this work we will use these new techniques as performed in  ATLAS:2020xyo . On the other hand, the tau jet with a displaced track provides an outstanding signature to our signal. Since the light pseudo-scalar is produced from the decay of a not-so-heavy Higgs boson ($\sim 125$ GeV ) in association with a rather heavy $Z$-boson ($\sim 91$ GeV), its momentum is usually not large, such that its decay products are usually soft. Considering the lower velocity of the taus from the pseudo-scalar decay, in this work we propose to reconstruct a merged (joint) vertex from the decays of the two taus in order to suppress the so-called pileup backgrounds. Although this reconstruction has larger uncertainties than the conventional (standard) vertexing, as will be demonstrated in our study, it can help to discriminate our signal from backgrounds.
@@ 2.2 Current theoretical and experimental constraints
In our scan $m_{12}^{2}$ is also a free parameter whose value is tuned to satisfy the theoretical constraints from the vacuum stability and perturbativity as well as unitarity. Because the Yukawa couplings of A are proportional to three different factors: $\kappa_{\ell}$ for leptons, $\kappa_{d}$ for down-type quarks, and $\kappa_{u}$ for up-type quarks, a light A will decay dominantly to $\tau^{+}\tau^{-}$ for $\kappa_{\ell}\gg\kappa_{u}$ and $\kappa_{\ell}\gg\kappa_{d}$. In addition, $\kappa_{u}$ and $\kappa_{d}$ can be stringently constrained by the searches for additional Higgs bosons via $gg\to H$, $gg\to H\bar{f}f$, $gg\to tbH^{-}$ and $gb\to tH^{-}$ production processes at the LHC. Therefore, for the Yukawa sector, we simply set $\kappa_{u}=\kappa_{d}$ to zero, and scan $\kappa_{\ell}$ within a range of small values,
@@
    The constraints by the signal strength of the 125 GeV Higgs. The non-vanishing $\cos(\beta-\alpha)$ in the couplings of $h$ to the SM particles, is constrained by HiggsSignals  Bechtle:2013xfa package to satisfy all the observations (within $2\sigma$). The decay $h\to AA$ is bounded by the invisible Higgs decay of $BR(h\to{\rm invisible})<10.7\%$  PhysRevD.110.030001 .
@@
i S,T,U
ii 125 GeV Higgs
iii additional Higgs
Figure 1: The respective exclusion capabilities of experimental constraints (i)-(iii).
@@
[…]
In Fig.2, we present scatter plots of the samples that satisfy all the aforementioned constraints on several other two-dimensional planes. The following characteristics can be observed from this figure:
@@
    The left panel shows that $\tan\beta$ is favored in the range 0.6 and 1.6, as explained in Appendix C.
@@
    The middle panel indicates that the invariant lifetime $c\tau$ of $A$ is primarily determined by $\kappa_{\ell}$ as expected and is inversely proportional to $\kappa_{\ell}^{2}$.
@@
    The right panel shows that the branching ratio of $h$ decay to $A$ and $Z$, i.e $BR(h\to ZA)$, is proportional to $\cos^{2}(\beta-\alpha)$, with a maximum value of approximately 0.068. Given that the cross-section for Higgs production via gluon fusion equals $48.58$ pb LHCHiggsCrossSectionWorkingGroup:2016ypw , the maximum cross section for $gg\to h\to ZA$ is approximately 3.3 pb.
@@ 3 Observability at the LHC
To save computing time in numerical computation, only a few representative benchmark points are chosen from the allowed parameter space for an illustration, as listed in Table 1. We choose low mass at 5 GeV and high mass at 30 GeV, and vary the invariant lifetime from 3 mm to 300 mm. We use FeynRules Alloul:2013bka to generate the UFO model file, MadGraph5_aMC@NLO Alwall:2014hca ; Frederix:2018nkq to generate parton-level events at the leading order (LO), Pythia8 Bierlich:2022pfr ; Sjostrand:2014zea for parton showering and hadronization, and then prepare signal events in HepMC3 format.
The gluon-gluon fusion Higgs production cross section is scaled to the N3LO value LHCHiggsCrossSectionWorkingGroup:2016ypw . For the $t\bar{t}$ background, we require the top quark to decay into a $W$-boson and a $b$ quark, with the $W$-boson subsequently decaying leptonically, while the $t\bar{t}$ cross section is scaled to the NNLO+NNLL value Czakon:2011xx . Among the $Z\tau\tau$ backgrounds, the $Z\gamma^{\ast}$ component is estimated to be about $38\%$ by MadGraph5_aMC@NLO. As a crosscheck, we show the distributions of the radius of particle production position in Fig.3, which demonstrates that these samples are generated according to physics expect.
[…]
@@ 3.1 The fast simulation and selection procedure
The framework Delphes deFavereau:2013fsa is adopted in this work. The setup of detector responses mainly follows that of ATLAS with pileup therein. New modules are developed as described below to accommodate a long-lived particle, in consistency with the methodology described in ATLAS:2023nze ; CMS:2022prd ; atlas:2019pn ; ATLAS:2019wqx ; CMS:2021tkn  [^3] These algorithms and the response parameters therein are followed with the working point Medium employed as default, just for a quick estimation of physics potential.:
@@
    Displaced tracks. In the scenario of LLP with a lifetime around dozens of millimeters, tracks from the LLP decay will be produced within the Tracker but somewhat far away from the collision point. Typically this can be reflected by the transverse impact parameter $d_{0}$ as shown in Fig.4. In our tracking module of Delphes, $d_{0}\geq 2.5$ mm is adopted to define the displaced tracks, while other setup as in ATLAS:2023nze . There is also an observation that the mass difference around 35 GeV $(\simeq m_{h}-m_{Z})$ leaves only less energy to tracks, so a low threshold $p_{T}\geq 0.5$ GeV is used for the track transverse momentum. But a track produced too far ($R_{v}\geq 300~{}mm$) will leave too few effective hits within the Inner Detector and thus will be rejected. From Figs.3 and 4 we see that with a fixed $c\tau$ value a lighter pseudo-scalar is more likely to travel a longer distance because it tends to have a higher velocity.
@@
    Displaced $\tau$ jet ($\tau$ jet with displaced tracks). A tau jet basing on the performance described in atlas:2019pn is expected to be associated with one or three extra displaced tracks within a cone $\sqrt{{(\Delta\phi)}^{2}+{(\Delta\eta)}^{2}}\leq 0.2$, as aforementioned. To keep as many tau signal events as possible, the threshold of tau $p_{T}$ is reduced to 10 GeV from the conventional cut like 20 GeV. For the taus with $p_{T}$ below 20 GeV, another (tight) working point atlas:2019pn ; CMS:2022prd is adopted for their identification, then the possibility for a b-jet to be mis-identified as a tau jet is controlled to about $10\%$. This tighter choice also brings an efficiency degradation for tau reconstruction roughly by $60\%$. For the events with only one hadronic tau tagged, a displaced muon is required as a substitution.
@@
    Reconstruction of $Z$-boson mass and visible mass of Higgs boson. As usual, the invariant mass for a pair of same-flavor and opposite-sign leptons is required to fall in a window of $81{\rm~{}GeV}<m_{\ell\ell}<101{\rm~{}GeV}$, with isolated leptons in the phase space of $p_{T}^{\ell}>15{\rm~{}GeV}$ and $|\eta|<2.5$. No doubt this $Z$-peak (in fact the leptons’ isolation) can suppress the QCD backgrounds down to a sufficiently low level. Also, by adding up the four-momenta of a pair of charged leptons (from $Z$) and a pair of tau candidates, a visible mass is reconstructed to suppress events other than the Higgs decay. As shown in Fig.6, it helps to reject all backgrounds. At last, the leptons from Z provide a very good trigger in actual experiments.
@@
Table 2: The separately estimated efficiencies for the backgrounds and the signals at each benchmark points. And the cross-sections prior to cuts of each backgrounds and signal (with the assumption $BR(h\to ZA)=1\%$ and $BR(A\to\tau\tau)=100\%$) are shown in the second column.
cross-section (pb) $m_{Z}$-window displaced $\tau$-jet displaced vertex $m^{vis}_{h}\leq 125$ GeV final efficiency
$Z\tau\tau$ 0.027 0.4 $1.4\times{10}^{-3}$ $7.9\times{10}^{-2}$ $7.5\times{10}^{-3}$ $3.3\times{10}^{-7}$
$Zbb$ 68.2 0.4 $9.8\times{10}^{-5}$ $6.9\times{10}^{-2}$ $9.9\times{10}^{-2}$ $4.1\times{10}^{-7}$
$t{\bar{t}}$ 92.4 $2.2\times{10}^{-2}$ $2.5\times{10}^{-4}$ $9.1\times{10}^{-2}$ $2.5\times{10}^{-2}$ $1.7\times{10}^{-8}$
BP1 0.014, assuming $BR(h\to$ $ZA)=1\%$, $BR(A\to$ $\tau\tau)=100\%$. 0.44 $3.2\times{10}^{-3}$ $8.2\times{10}^{-2}$ 0.95 $0.051\%$
@@
The performance of these simulation procedures is summarized in Table 2. For illustration, the efficiency is shown independently at each step. In fact, there are correlations among these cuts, especially between displaced tau-jet and displaced vertex, so the final efficiency differs from the simple product of these steps. In short, the signal events can be selected at a few thousandths, while the backgrounds are effectively (mainly by displaced tracks) reduced to about 4 or 5 orders lower. By a closer glance at these benchmark points, the efficiency for the high mass $(30\rm GeV)$ can reach the maximal value with $c\tau\simeq 100{\rm~{}mm}$ ($0.43\%$ for BP7). In comparison, the efficiencies for a lighter pseudo-scalar are relatively lower: it either contributes mainly softer taus, or goes beyond the sensitive zone since it travels faster (e.g. BP4). During these selection procedures, we find that the pileup entered by contaminating displaced tracks, playing a dominant role in backgrounds.
Our study can be applied to other models if they have the same event signatures. Upon these signatures, a future lepton collider like CEPC or FCCee will give better performance for the absence of pileup. But this goes beyond the scope of this work.
@@ 3.2 Results and discussions
Figure 7: $BR(h\to ZA)$ is presumed at ${10}^{-3}$ just for a schematic illustration of signal event yields, with the red for $(m_{A}=30{\rm~{}GeV},~{}c\tau=30{\rm~{}mm})$ and orange for $(m_{A}=5{\rm~{}GeV},~{}c\tau=30{\rm~{}mm})$. Other colors denote the backgrounds.
Table 3: Expected observation upon pure backgrounds for aforementioned benchmark points. The first row is the expected limit on $BR(h\to ZA)$ estimated with pure backgrounds, while the second row is the expected $BR(h\to ZA)$ to reach the expected $3\sigma$ significance through the selection procedures.
$(m_{A},c\tau)$ (5,3) (5,10) (5,30) (5,100) (30,10) (30,30) (30,100) (30,300)
[…]